OpenStudy (anonymous):

Gina pumps 40 cubic inches of air per second into a balloon. The function below shows the total amount of air f(t) inside the balloon after t seconds: f(w) = 40t + 5 Which statement best describes the y-intercept of the graph of the function? It is equal to 40 and represents the amount of air already present in the balloon before Gina starts blowing. It is equal to 5 and represents the amount of air already present in the balloon before Gina starts blowing. It is equal to 5 and represents the amount of air in the balloon after 1 second.

OpenStudy (anonymous):

It is equal to 45 and represents the amount of air in the balloon after 1 second.

OpenStudy (anonymous):

to find y-intercept you have have to make x = 0 or as per your function t = 0 and you find y = 5

OpenStudy (anonymous):

to the 4 other questions you do f(w) = 40t + 40 f(w) = 40t + 5 f(w) = 40t - 35 f(w) = 40t + 5 Check on your book.. the hint here is, you can change "40t" , as it says on the question...

OpenStudy (anonymous):

all you can do is change the constant

OpenStudy (anonymous):

so the answer would be what?

OpenStudy (anonymous):

yep I'll draw the function 1 min

OpenStudy (anonymous):

 your graph is something like this

OpenStudy (anonymous):

f(w) = 40t + 5 W = air in ballon t = time in seconds

OpenStudy (anonymous):

each function that I posted is the answer for each of your questions.. your Y-Intercept happen when t=0.. so f(w) = 40*(0) + 5 then f(w) = 5.. thats your y-intercept f(w) = 40t + 40 f(w) = 40t + 5 f(w) = 40t - 35 f(w) = 40t + 5